# TabiBERT:A Large-Scale ModernBERT Foundation Model and Unified Benchmarking Framework for Turkish

### v1

@@ A.1.5 Question Answering
As for XQuAD, limited volume of training data ($833$ training and only $179$ test examples) likely constrains model learning, resulting in lower scores across all models irrespective their context handling capabilities.
@@ A.1.6 Information Retrieval
We evaluate information retrieval performance on six datasets: BiText, MsMarco-TR, Scifact-TR, Fiqa-TR, NFCorpus-TR, and Quora-TR, which are translated into Turkish by the TRMTEB group (BaysanGungor2025TR-MTEB). Table 11 reports the NDCG@10 scores achieved by each model. Across the TRMTEB retrieval benchmarks, TabiBERT consistently outperforms existing Turkish encoder models, achieving an overall NDCG@10 score of $75.44$ and surpassing prior monolingual models such as BERTurk. We attribute these gains in part to TabiBERT’s exposure to more diverse pre-training data, including bilingual sources, which appear to be particularly beneficial for retrieval-oriented tasks. TabiBERT attains the highest NDCG@10 scores on the BiText and MsMarco-TR datasets. On Scifact-TR, TabiBERT trails the best-performing model, YTU-Cosmos-BERT, by approximately $6$ points (best score: $79.88$). However, Scifact-TR is the smallest dataset in the TRMTEB suite, containing only $339$ test instances, and is therefore the least representative benchmark in the retrieval evaluation. For Fiqa-TR, TabiBERT achieves the second-highest NDCG@10 score of $50.67$, exceeding BERTurk by $1.21$ points. Similarly, on NFCorpus-TR, TabiBERT ranks second with an NDCG@10 score of $31.19$, only $1.28$ points below the best-performing model, BERTurk. On Quora-TR, all models except TurkishBERTweet, including mmBERT, exhibit nearly identical performance, with scores clustered around $92.50$, and TabiBERT performing slightly below this level.
@@ A.1.7 Code Retrieval
We evaluated code retrieval performance using four datasets: Apps-TR, CosQA-TR, StackOverflowQA-TR, CodeSearchNet-21K-subset-TR (TabiBERTCodeRetrievalDatasets). Table 12 presents the NDCG@10 scores obtained by each model. In code retrieval tasks, TabiBERT outperforms all other Turkish models, achieving an average NDCG@10 score of $56.95$, considerably higher than the next best result of $54.54$ by BERTurk. Thanks to its support for longer contexts and the incorporation of code in pre-training data, the model demonstrates clear strengths in code retrieval tasks. Across the individual tasks, TabiBERT secures the top NDCG@10 score in three out of four datasets, with being the second highest in the Apps-TR, lagging behind only by $0.69$ points.
@@
A.1.8 Academic Domain Tasks
We evaluated academic domain tasks using four datasets: MedNLI-TR, PubmedRCT-10K-TR, SciCite-TR, ThesisAbstractClassification-11K (TabiBERTAcademicDomainDatasets). The MedNLI-TR dataset is an NLI dataset, the other three datasets are text classification datasets. Table 13 presents the results obtained by each model, reporting the macro F1 scores. We observe substantial improvements in academic domain tasks, such as NLI and multi-class text classification. For PubmedRCT-10K-TR task, TabiBERT shares the second place with YTU-Cosmos-BERT with a score of 75.32, only 0.29 behind the top performer BERTurk. On SciCite-TR, TabiBERT achieves a score of $83.29$, closely trailing YTU-Cosmos-BERT’s leading score of $84.09$ with a difference of $0.80$. On the demanding task of Thesis-Abstract-Classification-11K, characterized by long input contexts (with a 95th percentile context length of $1,152$ tokens) and sparse number of samples for each class (a sum of 60 examples per class throughout all splits), TabiBERT attains the highest score among Turkish models with $50.77$, outperforming the next best model BERTurk by $1.57$ points. Lastly, on the task of MedNLI-TR, TabiBERT achieves the highest score of $80.85$, slightly surpassing the second-best YTU-Cosmos-BERT by $0.23$ points. By leveraging a pre-training dataset rich in Turkish scientific articles and graduate theses, TabiBERT sets a new benchmark among Turkish models. These results highlight how TabiBERT’s exposure to academic domain texts and ability to process longer inputs underpin its strong performance in this area.
Table 13: Downstream performance of models on Academic domain tasks. The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.

### v3

@@ A.1.5 Question Answering
As for XQuAD, the limited volume of training data ($833$ training and only $179$ test examples) likely constrains model learning, resulting in lower scores across all models irrespective of their context handling capabilities.
@@ A.1.6 Information Retrieval
We evaluate information retrieval performance on six datasets: BiText, MsMarco-TR, Scifact-TR, Fiqa-TR, NFCorpus-TR, and Quora-TR, which are translated into Turkish by the TR-MTEB group (BaysanGungor2025TR-MTEB). Table 11 reports the NDCG@10 scores achieved by each model. Across the TRMTEB retrieval benchmarks, TabiBERT consistently outperforms existing Turkish encoder models, achieving an overall NDCG@10 score of $75.44$ and surpassing prior monolingual models such as BERTurk. We attribute these gains in part to TabiBERT’s exposure to more diverse pretraining data, including bilingual sources, which appear to be particularly beneficial for retrieval-oriented tasks. TabiBERT attains the highest NDCG@10 scores on the BiText and MsMarco-TR datasets. On Scifact-TR, TabiBERT trails the best-performing model, ytu-cosmos-BERT, by approximately $6$ points (best score: $79.88$). However, Scifact-TR is the smallest dataset in the TR-MTEB suite, containing only $339$ test instances, and is therefore the least representative benchmark in the retrieval evaluation. For Fiqa-TR, TabiBERT achieves the second-highest NDCG@10 score of $50.67$, exceeding BERTurk by $1.21$ points. Similarly, on NFCorpus-TR, TabiBERT ranks second with an NDCG@10 score of $31.19$, only $1.28$ points below the best-performing model, BERTurk. On Quora-TR, all models except TurkishBERTweet, including mmBERT, exhibit nearly identical performance, with scores clustered around $92.50$, and TabiBERT performing slightly below this level.
@@ A.1.7 Code Retrieval
We evaluated code retrieval performance using four datasets: Apps-TR, CosQA-TR, StackOverflowQA-TR, CodeSearchNet-21K-subset-TR (TabiBERTCodeRetrievalDatasets). Table 12 presents the NDCG@10 scores obtained by each model. In code retrieval tasks, TabiBERT outperforms all other Turkish models, achieving an average NDCG@10 score of $56.95$, considerably higher than the next best result of $54.54$ by BERTurk. Thanks to its support for longer contexts and the incorporation of code in pretraining data, the model demonstrates clear strengths in code retrieval tasks. Across the individual tasks, TabiBERT secures the top NDCG@10 score in three out of four datasets, with being the second highest in the Apps-TR, lagging behind only by $0.69$ points.
@@
A.1.8 Academic Understanding Tasks
We evaluated academic understanding tasks using four datasets: MedNLI-TR, PubmedRCT-10K-TR, SciCite-TR, ThesisAbstractClassification-11K (TabiBERTAcademicDomainDatasets). The MedNLI-TR dataset is an NLI dataset; the other three datasets are text classification datasets. Table 13 presents the results obtained by each model, reporting the macro F1 scores. We observe substantial improvements in academic understanding tasks, such as NLI and multi-class text classification. For PubmedRCT-10K-TR task, TabiBERT shares the second place with ytu-cosmos-BERT with a score of 75.32, only 0.29 behind the top performer BERTurk. On SciCite-TR, TabiBERT achieves a score of $83.29$, closely trailing ytu-cosmos-BERT’s leading score of $84.09$ with a difference of $0.80$. On the demanding task of Thesis-Abstract-Classification-11K, characterized by long input contexts (with a 95th percentile context length of $1,152$ tokens) and sparse number of samples for each class (a sum of 60 examples per class throughout all splits), TabiBERT attains the highest score among Turkish models with $50.77$, outperforming the next best model BERTurk by $1.57$ points. Lastly, on the task of MedNLI-TR, TabiBERT achieves the highest score of $80.85$, slightly surpassing the second-best ytu-cosmos-BERT by $0.23$ points. By leveraging a pretraining dataset rich in Turkish scientific articles and graduate theses, TabiBERT sets a new benchmark among Turkish models. These results highlight how TabiBERT’s exposure to academic understanding texts and ability to process longer inputs underpin its strong performance in this area.
Table 13: Downstream performance of models on Academic understanding tasks. The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.